abs l 1/2x +2l = l 3/4x -2 l Find the solution set for the equation
Answer by stanbon(57377)

abs l 1/2x +2l = l 3/4x -2 l Find the solution set for the equation
-------------------
(1/2)x +2 = (3/4)x-2 or (1/2)x + 2 = -(3/4)x+2
(1/4)x = 4 or (5/4)x = 0
x = 16 or x = 0
----------------------
Cheers,
Stan H.